# Optimization via Quantum Preconditioning

### v1

@@ IV.3.1 Compilation and Efficient Swap Network for Sparse Graphs
Superconducting quantum hardware have a given topology with a fixed connection between the qubits. This requires the use of two-qubit SWAP gates for achieving quantum logical operations between arbitrary non-neighboring qubits. Compiling QAOA circuits for sparse graph problems using a default swap network (i.e., on a linear arrangement of qubits [70, 78]) is far from efficient in terms of circuit depth and number of two-qubit gates. This sets fidelity requirements much higher than what is actually needed under the right compilation strategy. An efficient swap network [79, 80] for light-cone-induced subgraphs on random $3$-regular graph maximum-cut problems can reduce the number of ISWAP gates by about $52\%$, $73\%$, and, $85\%$ at $p=1$, $p=2$, and $p=3$, respectively [45]. Such smaller circuits would lead to much higher overall output fidelities. Therefore, the sampling requirements for reliably estimating expectation values would be much lower as well—in addition to shorter circuits, reducing the execution time—thus greatly speeding up the preconditioning.
For instance, a light-cone-induced subgraph has at most $N=43$ vertices at $p=3$ for a $3$-regular graph. Following Sec. IV.2.2, this translates in about $n_{\textrm{2Q}}\sim O(10^{4})$ via a default swap network or $n_{\textrm{2Q}}\sim O(10^{3})$ with an efficient compilation [79, 80, 45]. Assuming an average two-qubit gate fidelity $f=99.5\%$, the default strategy yields an overall fidelity of $F\simeq 10^{-22}$, and $F\simeq 10^{-2}$ for the efficient one. This highlights the importance of the compilation step on noisy quantum devices.
@@ IV.3.2 Decomposition Technique
We used the light-cone decomposition technique for classical emulations of QAOA circuits for sparse large $N$ problems, such as random $3$-regular graphs, grid energy problem (Sec. III.2.1). At shallow circuit depth, it trades a single $N$-qubit QAOA circuit for $O(N)$ QAOA circuits with $O(\exp p)$ qubits each. While this can be advantageous for small values of $p$, this requires the execution of many subcircuits. For instance, the preconditioning time via state vector emulations of the grid energy problem (Fig. 4) took $0.00270(2)$ s at $p=1$ and $32.9(2)$ s at $p=2$ on a 64Gb MacBook Pro with an Apple M1 Max chip. The $p=1$ and $p=2$ preconditioning times are within and out of budget, respectively. However, they do not account for the variational search of optimal QAOA angles. A long preconditioning time could be mitigated by parallelizing the execution of subcircuits on quantum hardware if more qubits than necessary are available (e.g., about twenty $50$-qubit circuits can be executed at a time on a $1,000$-qubit device).
Moreover, the light-cone decomposition technique is only applicable to sparse graphs. Dense graphs, such as the Sherrington-Kirkpatrick spin glasses considered cannot benefit from it. Because running the QAOA on large dense graph problems is out-of-scope for noisy devices due to the large number of qubits and gate requirements, this raises the question on whether other decomposition techniques could be envisioned in the context of quantum preconditioning [81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91].
@@ IV.3.3 Variational Optimization
For the random $3$-regular graph maximum-cut problems and Sherrington-Kirkpatrick spin glasses, we employed tabulated near-optimal parameters [29, 52, 27], thus skipping altogether the variational optimization. However, the uniqueness of the grid energy optimization problem required a variational optimization of the circuit parameters. We note that classes of optimization problems have been found to share near-optimal angles and that such a strategy of transferring parameters from one problem to another could be leveraged [94, 95, 29, 52, 96, 27, 28, 97, 57, 98, 99, 31]. Other strategies have also been advanced based on interpolation from lower $p$ values and quantum annealing scheduling [95, 100, 101]. However, this raises the question of the performance of the quantum preconditioner based on nonoptimal angles.
@@ V Conclusion
We introduced a quantum preconditioning method for quadratic unconstrained binary optimization problems. Here, preconditioning refers to the task of transforming the input problem into a more suitable form for a solver by improving its convergence. The proposed quantum preconditioner is based the QAOA [25, 26]. We demonstrated that state-of-the-art classical solvers such as simulated annealing and the Burer-Monteiro method benefited from working with the preconditioned input instead of the original one on a range of problems, including Sherrington-Kirkpatrick spin glasses, random $3$-regular graph maximum cut, and a real-world grid energy problem. The amount of preconditioning is controlled by the quantum circuit depth and renders the problem trivial for a classical solver in the infinite-depth limit. We showed that an increased circuit depth for shallow circuits can improve the convergence. Because classical optimization heuristics operate on an accuracy versus run-time basis, any additional time taken for preconditioning a problem should be given to the purely classical solution. Accounting for this, we showed that quantum preconditioning translates into a practical quantum-inspired advantage for random 3-regular graph maximum-cut problems through classical state vector emulations of the quantum circuits. Finally, we discussed the challenges and prospects for delivering a hardware-based quantum utility in optimization via quantum preconditioning.
@@
It would be interesting to investigate quantum preconditioning in the context of constrained problems—here we only considered unconstrained problems. Hard constraints are a part of many optimization problems. Yet, they lack methods for efficiently handling them on contemporary, noisy, quantum computers. Constraints can be encoded in quantum optimization in several ways, and each has their unique challenges. For example, constraints are often encoded as penalty terms [41], but bit strings may violate the penalties. In some cases, special mixers are used to keep the state within the constraint-satisfying subspace of the Hilbert space [102, 103], or the problem is encoded such that the Hilbert space contains only constraint-satisfying solutions [104, 105, 106, 107, 108, 109]. In the former (latter) case, the mixer (phase separator) requires multi-qubit gates. In our approach, one could quantum precondition a problem, in any formulation of the constraints in the problem, while letting a classical solver handle the hard constraints.
@@
Improving the fidelity of noisy hardware remains the backbone for quantum computers to potentially deliver a quantum advantage in the absence of quantum error correction. While it is common to leverage error mitigation techniques to enhance noisy expectation values [110, 111, 112], one should be mindful of the added time overhead of some of these techniques in a fair benchmark against classical optimization solvers. Error mitigation methods such as randomized compilation [113] or readout error mitigation [114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126] may provide a benefit with little overhead. Besides, exploring quantum preconditioning or the QAOA [127] within the framework of quantum error correction would also be interesting, especially with the advent of fault-tolerant demonstrations [128, 129, 130]. In particular, quantum error correction induces a run-time overhead, which then raises the question on whether there might be an opportunity for a quantum advantage for quantum preconditioning in a fault-tolerant era.
@@ Acknowledgements.
M.D. conceived and led the project with support from B.S. M.D. and T.O. performed simulations, data collection, and data analyses. M.D. wrote the manuscript with input from B.S. All co-authors contributed to the discussions leading to the completion of this project.
@@
The problem instance “South Carolina 500-Bus System: ACTIVSg500” formatted as a MEPS problem is publicly available at doi.org/10.5281/zenodo.14921060.
@@ A.1 Random $3$-Regular Graph Maximum-Cut Problems
Random $3$-regular graph maximum-cut problems are a standard testbed for investigating the performance of quantum optimization [25, 55, 52, 29, 131, 132, 133, 49, 134, 45, 135, 136, 137, 138, 139]. They have $N$ vertices, each connected by an edge to three other vertices at random. The edges carry unit weight. These graphs are sparse with the adjacency matrix $\mathsf{W}$ containing only three nonzero entries per row or column. Despite their simple apparent topology, it is NP-hard to devise an algorithm to find the maximum cut that can guarantee an approximation ratio of at least $99.7\%$ [140] with the current best-known classical algorithms yielding $\alpha\simeq 93.3\%$ [141]. We generated $3$-regular graph graphs at random using the Python package NetworkX [142, 143, 144].
The heuristics considered in this work may potentially deliver a solution with a higher approximation ratio but cannot guarantee it a priori. Random $3$-regular graph maximum-cut problems have been independently investigated on a range of solvers such as BM [145], extremal optimization [146], quantum annealing, and SA [136, 147, 148]—interestingly, it was shown that quantum annealing is less efficient than SA [147, 148]. Classical quantum-inspired algorithms based on simulated coherent Ising machines [149] and simulated bifurcation [150, 151] have also been considered [152]. Random $3$-regular graph maximum-cut problems have also served as a testbed for combinatorial optimization solvers based on graph neural networks [153, 154], although their performance is far from that of state-of-the-art methods [155, 156, 157]. We note that a one-to-one benchmark with the existing literature can be difficult as it focused on the time-to-solution (TTS) metric [158], which relates to the probability of finding the optimal solution. In this work, we measure the performance through the average approximation ratio instead, as quantum preconditioning cannot guarantee the same optimal solution between the original and preconditioned problems, except in the asymptotic $p\to+\infty$ limit. Besides, it might be interesting to investigate quantum preconditioning in the context of other state-of-the-art classical solvers than the BM solver and SA considered in this work, such as extremal optimization [146] and quantum-inspired algorithms [149, 158, 150, 151, 152].
@@ A.2 Sherrington-Kirkpatrick Spin Glasses
Sherrington-Kirkpatrick (SK) models [43] correspond to complete graphs, where the edge weights are independent and identically distributed random variables drawn from a normal distribution of zero mean and unit width $\mathsf{W}_{ij}\sim\mathcal{N}(0,1)$ with $\mathsf{W}_{ij}=\mathsf{W}_{ji}$. As a result, the adjacency matrix $\mathsf{W}$ encoding the problem belongs to the Gaussian orthogonal ensemble. Such graphs present two stark differences with $3$-regular graphs: First, they are as dense as possible because of the all-to-all connectivity between vertices. Second, they are weighted graphs.
It is NP-hard to find the optimal solution to an SK spin glass. However, an approximate solution arbitrary close to the optimal one can be obtained in polynomial time through a recently introduced approximate message passing algorithm [159]: A solution to an $N$-variable problem with an approximation ratio of at least $\alpha=(1-\varepsilon)$ for $\varepsilon>0$ can be found in $O(P_{\varepsilon}N^{2})$ with $P_{\varepsilon}$ a polynomial in $\varepsilon^{-1}$. Nevertheless, SK problems remain highly relevant for developing an understanding thanks to their long connection with statistical physics and benchmarking optimization solvers [53, 49, 89, 27, 54, 160, 90].
A.3 Power Grid Optimization Problem
Number of QAOA layers Optimized QAOA angles Objective value Number of nonzero terms
Original $226$
$p=1$ $\gamma_{1}\simeq-0.008418$ $\beta_{1}\simeq 2.757259$ $\simeq 59,721.27$ $600$
$p=2$ $\gamma_{1}\simeq-0.006541$ $\beta_{1}\simeq 1.074866$ $\gamma_{2}\simeq-0.012537$ $\beta_{2}\simeq 1.307313$ $\simeq 60,202.88$ $1,625$
@@ A.3 Power Grid Optimization Problem
Quantum preconditioning on a real-world power grid optimization problem seeking to compute the MPES of an energy network [57, 58, 59, 60] was investigated in Sec. III.3. We considered a realistic $500$-bus power system dataset “ACTIVSg500” mimicking the energy grid of the state of South Carolina in the United States of America [44]. The dataset was designed as part of the ARPA-E’s GRID DATA program to be similar to the actual electric grid while containing no confidential critical energy infrastructure information [61, 62, 63, 64, 65]. We show in Tab. 1 information regarding the QAOA settings that were used for preconditioning the $N=180$-variable problem. The reported QAOA angles correspond to minimizing the objective function $\tilde{C}_{\textrm{MPES}}(\boldsymbol{z})=\frac{1}{2}\sum_{i,j=1}^{N}\mathsf{W}_{ij}z_{i}z_{j}$. The reported objective value is that of the underlying maximum-cut problem of Eq. (6).
@@ Appendix B Quantum Preconditioning Eases Problems Hardness
Figure 7: (A) (B) $N=16$ random $3$-regular graph maximum cut problems. (C) (D) $N=16$ SK spin glasses. (A) (C) Average approximation ratio as a function of the number of classical iterations $n_{\textrm{iter}}$ for simulated annealing on various based on quantum preconditioning for various values of $p$. (B) (D) Average approximation ratio as a function of the number of QAOA layers $p$ used for quantum preconditioning for various number of iterations $n_{\textrm{iter}}$ in simulated annealing. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
Figure 8: (A) (C) $N$-variable random $3$-regular graph maximum cut problems. (B) (D) $N$-variable SK spin glasses. (A) (B) Average frustration index defined in Eq. (12) as a function of the number of QAOA layers $p$. (C) (D) Average normalized gap defined in Eq. (13) as a function of the number of QAOA layers $p$. Vertical dashed lines indicate the quantities evaluated on the original problem $\mathsf{W}$. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.

### v2

@@ IV.3.1 Compilation and efficient swap network for sparse graphs
Superconducting quantum hardware has a given topology, with a fixed connection between the qubits. This requires the use of two-qubit SWAP gates for achieving quantum logical operations between arbitrary non-neighboring qubits. Compiling QAOA circuits for sparse graph problems using a default swap network (i.e., on a linear arrangement of qubits [69, 77]) is far from efficient in terms of circuit depth and number of two-qubit gates. This sets fidelity requirements much higher than what is actually needed under the right compilation strategy. An efficient swap network [78, 79] for light-cone-induced subgraphs on random $3$-regular graph maximum-cut problems can reduce the number of ISWAP gates by about $52\%$, $73\%$, and, $85\%$ at $p=1$, $p=2$, and $p=3$, respectively [44]. Such smaller circuits would lead to much higher overall output fidelities. Therefore, the sampling requirements for reliably estimating expectation values would also be much lower—in addition to shorter circuits, reducing the execution time—thus greatly speeding up the preconditioning.
For instance, a light-cone-induced subgraph has at most $N=43$ vertices at $p=3$ for a $3$-regular graph. Following Sec. IV.2.2, this translates into about $n_{\textrm{2Q}}\sim O(10^{4})$ via a default swap network or $n_{\textrm{2Q}}\sim O(10^{3})$ with an efficient compilation [78, 79, 44]. Assuming an average two-qubit gate fidelity $f=99.5\%$, the default strategy yields an overall fidelity of $F\simeq 10^{-22}$, and $F\simeq 10^{-2}$ for the efficient one. This highlights the importance of the compilation step on noisy quantum devices.
@@ IV.3.2 Decomposition technique
We used the light-cone decomposition technique for classical emulations of QAOA circuits for sparse large-$N$ problems, such as random $3$-regular graphs and the grid energy problem (Sec. III.1.2). At shallow circuit depth, this trades a single $N$-qubit QAOA circuit for $O(N)$ QAOA circuits with $O(\exp p)$ qubits each. While this can be advantageous for small values of $p$, it requires the execution of many subcircuits. For instance, the preconditioning time via state-vector emulations of the grid energy problem (Fig. 4) took $0.00270(2)$ s at $p=1$ and $32.9(2)$ s at $p=2$ on a 64GB MacBook Pro with an Apple M1 Max chip. The $p=1$ and $p=2$ preconditioning times are within and out of budget, respectively; however, they do not account for the variational search of optimal QAOA angles. A long preconditioning time could be mitigated by parallelizing the execution of subcircuits on quantum hardware if more qubits than necessary are available (e.g., about twenty $50$-qubit circuits can be executed at a time on a $1,000$-qubit device).
Moreover, the light-cone decomposition technique is only applicable to sparse graphs. Dense graphs, such as the Sherrington-Kirkpatrick spin glasses considered, cannot benefit from it. Because running the QAOA on large, dense graph problems is out of scope for noisy devices due to the large number of qubits and gate requirements, this raises the question of whether other decomposition techniques could be envisioned in the context of quantum preconditioning [80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90].
@@ IV.3.3 Variational optimization
For the random $3$-regular graph maximum-cut problems and Sherrington-Kirkpatrick spin glasses, we employed tabulated near-optimal parameters [29, 51, 27], thus skipping altogether the variational optimization; however, the uniqueness of the grid energy optimization problem required a variational optimization of the circuit parameters. We note that classes of optimization problems have been found to share near-optimal angles and that such a strategy of transferring parameters from one problem to another could be leveraged [93, 94, 29, 51, 95, 27, 28, 96, 56, 97, 98, 31]. Other strategies have also been advanced based on interpolation from lower $p$ values and quantum annealing scheduling [94, 99, 100]; however, this raises the question of the performance of the quantum preconditioner based on nonoptimal angles.
@@ V Conclusion
We have introduced a quantum-preconditioning method for quadratic unconstrained binary optimization problems. Here, preconditioning refers to the task of transforming the input problem into a more suitable form for a solver by improving its convergence. The proposed quantum preconditioner is based the QAOA [25, 26]. We demonstrated that state-of-the-art classical solvers such as simulated annealing and the Burer-Monteiro method benefited from working with the preconditioned input instead of the original one on a range of problems, including Sherrington-Kirkpatrick spin glasses, random $3$-regular graph maximum cut, and a real-world grid energy problem. The amount of preconditioning is controlled by the quantum circuit depth and renders the problem trivial for a classical solver in the infinite-depth limit. We showed that an increased circuit depth for shallow circuits can improve the convergence. Because classical optimization heuristics operate on an accuracy versus run-time basis, any additional time taken for preconditioning a problem should be given to the purely classical solution. Accounting for this, we showed that quantum preconditioning translates into a practical quantum-inspired advantage for random 3-regular graph maximum-cut problems through classical state-vector emulations of the quantum circuits. Finally, we discussed the challenges and prospects for delivering a hardware-based quantum utility in optimization via quantum preconditioning.
The results call for developing a better understanding of what makes a good quantum preconditioner. We found that quantum preconditioning via the QAOA reduces the problem’s frustration, making local moves more efficient in approaching the global optimum. In this work, the use of the QAOA as a preconditioner was heuristic, and it is possible that other quantum circuits might be better suited.
@@
It would be interesting to investigate quantum preconditioning in the context of constrained problems—here, we only considered unconstrained problems. Hard constraints are a part of many optimization problems; however, there is a lack of methods for efficiently handling them on contemporary, noisy, quantum computers. Constraints can be encoded in quantum optimization in several ways, and each has its unique challenges. For example, constraints are often encoded as penalty terms [41], but bit strings may violate the penalties. In some cases, special mixers are used to keep the state within the constraint-satisfying subspace of the Hilbert space [101, 102], or the problem is encoded such that the Hilbert space contains only constraint-satisfying solutions [103, 104, 105, 106, 107, 108]. In the former (latter) case, the mixer (phase separator) requires multiqubit gates. In our approach, one could quantum precondition a problem in any formulation of the constraints while letting a classical solver handle the hard constraints.
@@
Improving the fidelity of noisy hardware remains the backbone for quantum computers to potentially deliver a quantum advantage in the absence of quantum error correction. While it is common to leverage error-mitigation techniques to enhance noisy expectation values [109, 110, 111], one should be mindful of the added time overhead of some of these techniques in a fair benchmark against classical optimization solvers. Error-mitigation methods such as randomized compilation [112] or readout-error mitigation [113, 114, 115, 116, 117, 118, 119, 120, 121, 70, 122, 123] may provide a benefit with little overhead. Furthermore, exploring quantum preconditioning or the QAOA [124] within the framework of quantum error correction would also be interesting, especially with the advent of fault-tolerant demonstrations [125, 126, 127]. In particular, quantum error correction induces a run-time overhead, which then raises the question of whether there might be an opportunity for a quantum advantage for quantum preconditioning in a fault-tolerant era.
@@ Acknowledgements.
M.D. conceived and led the project with support from B.S. M.D. and T.O. performed simulations, data collection, and data analyses. M.D. wrote the manuscript with input from B.S. All coauthors contributed to the discussions leading to the completion of this project.
@@
The problem instance “South Carolina 500-Bus System: ACTIVSg500” formatted as a MEPS problem is publicly available at doi.org/10.5281/zenodo.14921060 [128].
@@ A.1 Random $3$-regular graph maximum-cut problems
Random $3$-regular graph maximum-cut problems are a standard testbed for investigating the performance of quantum optimization [25, 54, 51, 29, 129, 130, 131, 48, 132, 44, 133, 134, 135, 136, 137]. They have $N$ vertices, each connected by an edge to three other vertices at random. The edges carry unit weight. These graphs are sparse, with the adjacency matrix $\mathsf{W}$ containing only three nonzero entries per row or column. Despite their apparently simple topology, it is NP-hard to devise an algorithm to find the maximum cut that can guarantee an approximation ratio of at least $99.7\%$ [138], with the current best-known classical algorithms yielding $\alpha\simeq 93.3\%$ [139]. We generated random $3$-regular graphs using the Python package NetworkX [140, 141, 142].
The heuristics considered in this work have the potential to deliver a solution with a higher approximation ratio, but they cannot guarantee it a priori. Random $3$-regular graph maximum-cut problems have been independently investigated on a range of solvers such as BM [143], extremal optimization [144], quantum annealing, and SA [134, 145, 146]—interestingly, it was shown that quantum annealing is less efficient than SA [145, 146]. Classical quantum-inspired algorithms based on simulated coherent Ising machines [147] and simulated bifurcation [148, 149] have also been considered [150]. Random $3$-regular graph maximum-cut problems have also served as a testbed for combinatorial optimization solvers based on graph neural networks [151, 152], although their performance is far from that of state-of-the-art methods [153, 154, 155]. We note that a one-to-one benchmark with the existing literature can be difficult as it has focused on the time-to-solution (TTS) metric [156], which relates to the probability of finding the optimal solution. In this work, we measure the performance through the average approximation ratio instead, as quantum preconditioning cannot guarantee the same optimal solution between the original and preconditioned problems, except in the asymptotic $p\to+\infty$ limit. Furthermore, it might be interesting to investigate quantum preconditioning in the context of state-of-the-art classical solvers other than the BM solver and SA considered in this work, such as extremal optimization [144] and quantum-inspired algorithms [147, 156, 148, 149, 150].
@@ A.2 Sherrington-Kirkpatrick spin glasses
SK models [42] correspond to complete graphs, where the edge weights are independent and identically distributed random variables drawn from a normal distribution of zero mean and unit width $\mathsf{W}_{ij}\sim\mathcal{N}(0,1)$ with $\mathsf{W}_{ij}=\mathsf{W}_{ji}$. As a result, the adjacency matrix $\mathsf{W}$ encoding the problem belongs to the Gaussian orthogonal ensemble. Such graphs present two stark differences with $3$-regular graphs: first, they are as dense as possible because of the all-to-all connectivity between vertices; second, they are weighted graphs.
It is NP-hard to find the optimal solution to an SK spin glass; however, an approximate solution arbitrarily close to the optimal one can be obtained in polynomial time through a recently introduced approximate message-passing algorithm [157]: a solution to an $N$-variable problem with an approximation ratio of at least $\alpha=(1-\varepsilon)$ for $\varepsilon>0$ can be found in $O(P_{\varepsilon}N^{2})$ with $P_{\varepsilon}$ a polynomial in $\varepsilon^{-1}$. Nevertheless, SK problems remain highly relevant for developing an understanding thanks to their long-standing connection with statistical physics and benchmarking optimization solvers [52, 48, 88, 27, 53, 158, 89].
A.3 Power-grid optimization problem
Number of
QAOA layers
Optimized
QAOA angles
Objective
value
Number of
nonzero terms
Original
$226$
$p=1$
$\gamma_{1}\simeq-0.008418$
$\beta_{1}\simeq 2.757259$
$\simeq 59,721.27$
$600$
$p=2$
$\gamma_{1}\simeq-0.006541$
$\beta_{1}\simeq 1.074866$
$\gamma_{2}\simeq-0.012537$
$\beta_{2}\simeq 1.307313$
$\simeq 60,202.88$
$1,625$
@@ A.3 Power-grid optimization problem
Quantum preconditioning on a real-world power-grid optimization problem seeking to compute the MPES of an energy network [56, 57, 58, 59] was investigated in Sec. III.3. We considered a realistic $500$-bus power system dataset “ACTIVSg500” mimicking the energy grid of the state of South Carolina in the United States of America [43]. The dataset was designed as part of ARPA-E’s GRID DATA program to be similar to the actual electric grid while containing no confidential information about critical energy infrastructure [60, 61, 62, 63, 64]. In Tab. 1, we present information regarding the QAOA settings that were used for preconditioning the $N=180$-variable problem. The reported QAOA angles correspond to minimizing the objective function $\tilde{C}_{\textrm{MPES}}(\boldsymbol{z})=\frac{1}{2}\sum_{i,j=1}^{N}\mathsf{W}_{ij}z_{i}z_{j}$. The reported objective value is that of the underlying maximum-cut problem of Eq. (6).
@@ Appendix B Quantum Preconditioning Eases Problems Hardness
Figure 7: (A) (B) $N=16$ random $3$-regular graph maximum-cut problems. (C) (D) $N=16$ SK spin glasses. (A) (C) Average approximation ratio as a function of the number of classical iterations $n_{\textrm{iter}}$ for simulated annealing on various, based on quantum preconditioning, for various values of $p$. (B) (D) Average approximation ratio as a function of the number of QAOA layers $p$ used for quantum preconditioning for various number of iterations $n_{\textrm{iter}}$ in simulated annealing. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.
Figure 8: (A) (C) NN-variable random $3$-regular graph maximum-cut problems. (B) (D) NN-variable SK spin glasses. (A) (B) Average frustration index defined in Eq. (12) as a function of the number of QAOA layers pp. (C) (D) Average normalized gap defined in Eq. (13) as a function of the number of QAOA layers pp. Vertical dashed lines indicate the quantities evaluated on the original problem $\mathsf{W}$. Each data point is averaged over $200$ randomly generated problem instances. Error bars indicate the standard error of the mean.